### v1

# Not flexible enough? Impacts of electric carsharing on a power sector with variable renewables11footnotemark: 1

Abstract
Electrifying the car fleet is a major strategy for mitigating greenhouse gas emissions in the transport sector. However, electrification alone will not solve all the negative externalities associated with cars. In light of other problems such as congestion, noise and street space as well as concerns about the use of mineral resources for battery electric cars, reducing the size of the car fleet would be beneficial, particularly in cities. Carsharing could offer a way to reconcile current car usage habits with a reduction in the size of the car fleet. However, it could also reduce the potential of electric cars to align their grid interactions with variable renewable electricity generation. We investigate how electric carsharing may impact the power sector in the future. To do so, we combine three open-source quantitative methods, including sequence clustering of car travel diaries, a probabilistic tool to generate synthetic electric vehicle time series, and an optimization model of the power sector. For 2030 scenarios of Germany with a renewable share of at least 80%, we show that switching to electric carsharing only moderately increases power sector costs. In our main setting with a mixed fleet including shared and private electric vehicles, carsharing increases yearly power sector costs by less than 100 euros per substituted private electric car. This cost effect increases with the share of shared cars in the overall fleet and is largest under the assumption of bidirectional charging. It is mitigated when other sources of flexibility for the power sector are considered, such as a flexible green hydrogen industrial demand. Carsharing further causes a shift from wind power to solar PV in the optimal capacity mix, and may also trigger additional investments in stationary electricity storage, especially in flexibility-constrained settings. Overall, we find that shared electric cars still have the potential to be operated largely in line with variable renewable electricity generation. We conclude that electric carsharing is unlikely to cause much damage to the power sector, but could bring various other benefits, which may outweigh power sector cost increases.
keywords:
Carsharing, Battery electric vehicles, Power sector flexibility, Sequence clustering
1 Introduction
Electrifying the passenger vehicle fleet is a major strategy for reducing greenhouse gas emissions in the passenger transport sector (noauthor_global_2023 (19, 4)). This is because battery electric vehicles (BEVs) often have lower life-cycle emissions than internal combustion engine vehicles (ICEVs) already today, and potentially much more so in the future (sacchi_when_2022 (34, 5, 22, 18)). As policymakers increasingly support the electric vehicle industry, the electrification of passenger cars is also becoming more and more widespread. This trend is expected to continue and accelerate as bans on the sale of new non-zero emission cars are announced in various parts of the world. In the European Union, the ban will take effect from 2035. This will also be the case in some states of the United States such as California. In China, at least half of all new cars sold after 2035 will have to be electric, fuel cell or plug-in hybrid.
At the same time, there is a growing interest in carsharing, especially in urban areas (munzel2020explaining (32, 12, 23)). By reducing the overall fleet size, carsharing has the potential to increase available street space in cities (creutzig2020fair (6, 14)). Carsharing and car electrification are interrelated, as the former could also foster consumer uptake of electric cars. Shared cars could thus play a role in bringing the electrification of the car fleet to scale (mccollum2018interaction (28)). However, while electric carsharing offers numerous benefits, it also poses some challenges for the power sector. One challenge is the potential loss of flexibility. Privately-owned BEVs with system-oriented charging capabilities can provide flexibility to the power sector. This facilitates the integration of variable renewable energy sources such as solar and wind power (liu2024 (24, 44, 41, 17)), and can help to shave peak loads (xu2018planning (45)). This way, BEVs can also reduce stationary electricity storage needs (xu2023storage (43)) and respective use of mineral resources (deetman2021 (8)). Electric carsharing may reduce the flexibility potential of BEVs due to their increased driving frequency, which translates into a lower grid availability, as well as lower aggregate battery capacity of the fleet.
Taking into account both the trends of electrification of passenger transport on the one hand and the development of sharing options on the other, we investigate the impacts of a switch from private BEVs to electric carsharing on a power sector dominated by variable renewable energy sources. In our study, carsharing refers to a free-floating fleet of shared vehicles that can be rented on a per-trip basis. Carsharing differs from ride-sharing or share-pooling in that individuals do not share a vehicle during a journey. In Germany, carsharing has been in use for longer than ride-pooling and is therefore more established (burghard2022sharing (3)). Carsharing might contribute to reducing car ownership (jochem_does_2020 (20, 31)). Therefore, carsharing can also be relevant in the transport decarbonization strategy, as a reduction in the size of the car fleet may be desirable beyond the mere electrification of the passenger road transport sector in order to achieve decarbonization targets (dirnaichner_life-cycle_2022 (10, 29)). Case studies have shown that carsharing can also help to reduce other negative externalities, and, for instance, increase the availability of public space in central city areas (diana_spatial_2022 (9)). Finally, carsharing could help mitigate negative impacts of cars on marine and freshwater ecotoxicity as well as on mineral resource scarcity (vilaca_life_2022 (42)). Hence, electric carsharing could also help tackle the potential trade-off between decarbonizing transport and the need for critical metals (zhang_trade-off_2023 (47, 46)).
In addition to the challenge of electrifying the passenger car fleet, meeting decarbonization targets also requires an increasing use of electricity generated from renewable energy sources. In most countries, the renewable energy sources with the largest potential are wind and solar power, which have variable generation patterns (lopezprol2021 (26)). This requires future power sectors to be more flexible. Various options for temporal and spatial flexibility exist, including various types of energy storage (schill2020 (35)). While electrifying passenger road transport increases electricity demand, and thus the need for integrating variable renewables, it may also increase the power sector flexibility if BEVs are charged smartly or bidirectionally using vehicle-to-grid (V2G) technologies (syla2024 (40, 27, 36, 25)). In particular, it has been found that BEVs may help balance daily variations of solar power (brown2018synergies (2)). One important factor enabling this flexibility is the fact that many privately-owned vehicles remain idle for extended periods of time. Assuming sufficient grid connection, electric vehicles could charge during times of abundant renewable power and, if bidirectional charging is possible, feed electricity back into the grid during times of scarcity. The actual flexibility potential that BEV can provide to the power sector depends on various factors, such as driving profiles and plug-in behaviours (gschwendtner_impact_2023 (16)). Furthermore, the overall battery capacity of the BEV fleet that is equipped with smart or bidirectional charging is also a significant factor.
A few analyses focus specifically on the potential of carsharing fleets for grid congestion management (brinkel2022grid (1)) or take the perspective of operators of shared BEV fleets (kahlen2018electric (21)) or the end users’ perspective to investigate whether carsharing users are willing to offer temporal flexibility when using shared BEVs (suel2024vehicle (39)). We aim to address this research gap with a quantitative, model-based study of the power sector effects related to switching from privately-owned BEVs to electric carsharing. We do so for 2030 scenarios of the German power sector with a renewable energy share of at least 80%. We use a combination of three open-source quantitative methods: (i) sequence clustering techniques on car travel diaries from a national German mobility survey, (ii) emobpy, a probabilistic tool for generating synthetic BEV time series, and (iii) DIETER, a capacity expansion model of the power sector. For clarity, we assume no changes in demand for car mobility, and consider three different BEV charging strategies which apply to the whole vehicle fleet.
2 A combination of three quantitative methods
Our methodological approach is based on a three-step procedure (Figure 1). First, probability distributions for car mobility are computed based on empirical data. These distributions are generated for both private and shared BEVs, assuming a constant substitution rate. Secondly, the open-source tool emobpy is used to generate time series for the electric consumption and charging availability of BEVs in an hourly resolution. Thirdly, these time series serve as input for a linear optimization open-source power sector model called DIETER. The first two steps are exclusively carried out using Python, while the third step also involves GAMS programming. All codes are available on the dedicated GitLab repository. [^2] The repository can be accessed at https://gitlab.com/diw-evu/electric-carsharing.
Figure 1: Flowchart of methods, main assumptions, inputs and outputs
2.1 Sequence clustering
To generate mobility probability distributions, we use travel diaries from the representative survey Mobilität in Deutschland (MiD) (eggs2018mobilitat (11)). The travel diaries are reshuffled into a sequence format, and a clustering algorithm is applied to identify distinct groups of car usage.
2.1.1 Travel diaries
GPS tracking and big data methods can provide access to highly detailed time series with high geographical and temporal resolution for existing car users. However, electric cars and carsharing are still in the early stages of development. Therefore, generalizing the behaviour of early adopters to a country-wide setting could lead to flawed conclusions. This study departs from existing empirical data on electric car or carsharing use and instead uses travel diaries from a representative survey, which allows to consider geographical components that are relevant in this context, such as the level of urbanization. In addition, travel diaries provide detailed information on car use at a high temporal resolution.
This study uses the B.1 version of the 2017 edition of the MiD, commissioned by the German Federal Ministry for Transport and Digital Infrastructure. [^3] Bundesministerium für Digitales und Verkehr (BMDV). The data collection took place in Germany from May 2016 to September 2017. A total of 316,361 individuals from 156,420 households were surveyed, representing the German population at the time (around 41 million households and 82 million individuals). For the analysis, we only considered car trips with available departure and arrival times. We excluded passenger trips to avoid double counting, as most car trips taken as a passenger are with another household member as the driver. We also excluded car trips undertaken during one’s professional activity. Individuals whose diary implied overlapping trips, i.e. at least two trips taking place at least partially at the same time, and trips that span two days were also removed. The final database comprises 355,136 car trips.
2.1.2 Sequencing
Travel diaries contain a wealth of information, including the timing of departure and arrival, trip duration, distance, and destination. Additionally, the serial aspect of the data can provide further insight into individuals’ car use. The sequence format allows for a large amount of information to be encapsulated into a one-dimensional object, making comparisons across individuals easier without losing any information. Hence, we recode travel diaries into sequences, which results in a sample of 119,519 sequences.
Sequence formatting requires two essential features: a fixed time interval and a taxonomy of mutually exclusive possible states. The time interval decomposes a given time frame into blocks of equal duration. In this case, the time frame is a full day from midnight until 23:59 o’clock. The time interval is five minutes, meaning that each day is broken into 288 blocks. By “sequence”, we refer to the succession of these 288 blocks that map the clock time. The following step in sequence formatting involves assigning each block to a single “state”. These states correspond to the variable of interest, which in this case is car mobility. We define two states: “idle” when the car is not moving and “on the move” when a car trip takes place.
2.1.3 Clustering
We implement the Hierarchical Agglomerative Clustering (HAC) algorithm using the Levenshtein distance as a distance measure and Ward’s method as the linkage criterion. The clustering is performed on subdatasets that partition the overall sequence dataset by location type: metropolis, big city, central city/medium-sized city, urban area, and small town/village area. [^4] The location typology corresponds to the variable RegioStatR Gem5 provided in the MiD dataset. Additional information is available on the BMDV website at this link. The final number of clusters is selected using dendrograms (Figure 7).
The computed descriptive statistics for the clusters (Table 3) indicate that car usage can be characterised by the daily distance travelled by car, which correlates well with the average trip distance and duration. For example, individuals in cluster 6 travel approximately 11 kilometers by car, with an average trip distance of 5.5 kilometers and an average trip duration of 14 minutes. In contrast, individuals in cluster 4 travel an average of 50 kilometers, with an average trip distance of 18.5 kilometers and an average trip duration of 34.2 minutes. To visualize differences in car use, we refer to sequence index plots (Figure 8).
This applies to all locations, even though there may be some differences in the characteristics of the clusters across locations. For example, in the cluster of users who drive short distances overall during the day, the total distance travelled as well as the average trip distance are greater in less urbanized areas. Conversely, the average trip duration is bigger in more urbanized areas, which is in line with congestion being more prevalent in big cities than rural areas.
2.2 Synthetic BEV time series
Using the open-source tool emobpy, [^5] For more detailed information on emobpy, we refer to the dedicated publication (gaete-morales_open_2021 (13)) and the appendix (B). we generate synthetic BEV time series for private and shared BEVs based on clustering results. Synthetic time series are computed at the cluster level for each location at a 5-minute resolution. In all computations, we use the default values for the Volkswagen ID.3 model. For shared BEVs across all clusters and private BEVs from the cluster with longest trips, we assume a battery capacity of 100 kWh, compared to 58 kWh in other cases. Probability distributions for charging station availability can be found in Table 5. The charging strategy used to compute the grid electricity time series is balanced.
2.2.1 Vehicle mobility
To generate the vehicle mobility time series, several input probability distributions have to be provided: the distribution of the number of trips per day, the distribution of the departure time of the trips and the joint distribution of trip duration and trip distance. We introduce new features to emobpy in order to generate time series based on more detailed distributions. More precisely, the distribution of the departure time is conditional on the total number of trips and the trip rank, on top of the trip destination and the day type. Similarly, the joint probability of duration and distance is also conditional on the trip destination, the day type and, as a new feature, the total number of trips. The conditionality on the total number of trips and the trip rank, where applicable, is only active for the generation of time series for private cars and not shared ones.
Shared BEVs
Assuming a substitution rate of five between private cars and shared cars, we derive probability distributions for shared BEVs based on the probability distributions for private BEVs. The distribution of number of trips for shared BEVs is the same as for private BEVs, but centered around a new mean (see B). We assume that the distribution of trip departure times remains the same as in the private case, except that we condition only on the trip destination. For the joint distribution of trip duration and distance, we also assume the same as in the private case and remove the conditioning on the total number of trips (Table 4). Thus, this modelling approach is based on the assumption that carsharing does not induce a change in the number or type of car trips made.
Weekly time series
Time series are computed at the cluster level for each location type on weekdays. On Saturdays and Sundays, time series are generated at the location level, as summarised in Table 4. This approach avoids making additional assumptions on how clusters should be matched across day types. One of our working hypotheses is thus that car usage during the week drives the likelihood of switching to carsharing, regardless of weekend usage.
2.3 Power system modelling
We apply the open-source power sector model DIETER (Dispatch and Investment Evaluation Tool with Endogenous Renewables). As a capacity expansion model, it minimizes fixed and variable power sector costs under various feasibility and policy constraints. The model is solved for a full year in an hourly resolution (8760 consecutive hours). Its main focus is the integration of variable renewable energy sources (vRES) by means of various options for power sector flexibility. We apply the model to a 2030 scenario of Germany with an 80% renewable share. For clarity and generalizability, we abstract from the European interconnection and model Germany as an electric island, assuming perfect expansion of the transmission and distribution grids within the country. Sector coupling with the heating sector is modelled assuming a stock of 7.5 million heat pumps, which are assumed to be operated inflexibly, i.e. to provide heat at the time of demand. [^6] For more details, see roth2023flexible (33). Using around 52 TWh of electricity, these heat pumps generate around 143 TWh of heat, corresponding to around twenty percent of the annual space heating demand from residential buildings in Germany. In additional model runs, we also include a 30 TWh flat annual demand for green hydrogen, which corresponds to an additional electricity demand of around 42 TWh.
Technologies and renewable energy constraint
[…]
Modeling interactions with BEVs
Sector coupling with BEVs is modelled using distinct vehicle profiles. These profiles represent individual vehicles which are scaled up to the overall fleet. In the case of uncontrolled charging, each profile is considered an exogenous electricity demand that is added to the overall demand. When considering smart or bidirectional charging, the model chooses when and how much electricity should flow into a vehicle, while ensuring battery levels that are compatible with the mobility constraint. This relies on the exogenous time series “driving electricity consumption” and “grid availability” derived with emobpy. Discharging electricity from BEV back to the grid incurs additional variable costs of 5 euros/MWh to account for battery degradation.
The model is calibrated based on a 15 million BEV fleet in 2030, which aligns with the target of the German coalition. The distribution of cars across locations is assumed to be the same as during the survey (Table 6). Additionally, it is assumed that clusters with very long trips (clusters 1 and 2 in each location) are not regular daily patterns, but rather trips undertaken under exceptional circumstances, such as holidays. As these patterns represent a very small proportion of overall sequences (around 2.4%), we exclude them from our analysis and assume that the 15 million BEV fleet is distributed across locations and clusters 3 to 6 (Table 1).
Metropolises Big cities Middle-size cities Small cities Rural areas Total
Cluster 3 120,308 82,729 157,940 295,736 337,853 994,566
Cluster 4 455,816 308,030 393,170 1,010,433 560,125 2,727,574
Cluster 5 486,317 517,490 1,112,303 907,747 1,364,750 4,388,607
Cluster 6 559,180 818,479 1,639,890 1,840,138 2,031,566 6,889,253
Total 1,621,621 1,726,728 3,303,303 4,054,054 4,294,294 15,000,000
Table 1: Assumed BEV distribution across locations and clusters
2.4 Scenarios
We investigate a number of scenarios which differ with respect to the size of the overall vehicle fleet, the number of shared cars, and the charging strategy. Scenarios are compared to respective reference settings without carsharing, i.e., references only include privately-owned BEV profiles.
Scenario families
We group scenarios into two broad families. The first family only considers BEVs that switch to carsharing in the scenario (“Shared only”). The second family additionally includes other electric cars that stay private both in the reference and the scenario (“Shared + other BEVs”). Hence, the BEV fleets considered in the latter family are way larger than the ones in the former (Table 2). The former family highlights the power sector effects of switching to carsharing, assuming away many other potential sources sources of flexibility. While results are instructive, they may not be policy-relevant. The latter family is designed to mimic a more realistic setting with additional BEVs that do not switch to car sharing, but remain privately-owned. Depending on the charging strategy, these additional “non-switching” BEVs also serve as a proxy for other flexibility options that are not included in the model. We thus consider “Shared + other BEVs” to be the main scenario family in the following.
Shared only Shared + other BEVs
Low uptake High uptake Low uptake High uptake
Reference Scenario Reference Scenario Reference Scenario Reference Scenario
Number of BEVs (in million) 4.9 1.0 7.9 1.6 15.0 11.1 15.0 8.8
Number of private BEVs (in million) 4.9 0.0 7.9 0.0 15.0 10.2 15.0 7.2
Number of shared BEVs (in million) 0.0 1.0 0.0 1.6 0.0 1.0 0.0 1.6
Share of shared BEVs (in %) 0.0 100 0.0 100 0.0 8.7 0.0 17.9
Number of substituted BEVs (in million) - 3.9 - 6.2 - 3.9 - 6.2
Overall BEV battery capacity (in GWh) 282 98 461 161 912 729 912 613
Substituted BEV battery capacity (in GWh) - 183 - 300 - 183 - 299
Overall EV power consumption (in TWh) 7.7 7.7 17.7 17.7 46.7 46.7 46.7 46.7
Table 2: BEV units, battery capacity and power consumption for all reference settings and scenarios
Carsharing uptake
For each scenario family, we consider two possible carsharing uptake regimes: a low and a high one. In the low uptake regime, only BEVs belonging to the cluster of shortest trips (cluster 6) are assumed to switch to carsharing. In the high uptake regime, we assume that all BEVs in metropolises switch to carsharing as well as most BEVs in big cities and some in the middle-size cities. In small cities, only cluster 6 is assumed to switch to carsharing. For both uptake regimes, we assume that there is no switch to carsharing in rural areas. The number of substituted private BEVs is approximately 3.9 million in the low uptake scenarios compared to 6.2 million in the high uptake scenarios. Further details are available in Figure 2 and Table 2. The low uptake regime is more conservative, as only users that rely the less on their cars switch to carsharing. The high uptake regime is more progressive and aims at illustrating a very optimistic scenario when analyzing impacts on the power system. Average grid availability for each hour of the week is illustrated in Figure 9.
Figure 2: Definition of carsharing uptake regimes based on cluster and location types. Note that this represents the reference fleet of the “Shared + other BEVs” scenario family. For the “Shared only” family, the reference fleet is only constituted of the shaded areas that switch to shared BEVs in the scenario.
Charging strategy
For each scenario, we consider various charging strategies: uncontrolled charging, smart charging, or bidirectional charging. To not mix up effects, we assume that both the reference and the respective scenarios always have the same charging strategy. Additionally, only one charging strategy is used for the entire fleet at a time to enable meaningful comparisons. This means that all BEVs charge smartly in the reference and the scenarios, or they all charge bidirectionally, etc. We consider different charging strategies since it is unclear which charging strategy will be mostly used by BEVs in the future. While it is unlikely that the whole BEV fleet will be operated in an uncontrolled manner (schill2015power (36)), it also seems unlikely that the whole fleet will be operated bidirectionally. By considering these strategies and homogeneous fleets, we analyze extreme settings which can be thought of as boundaries for outcomes of alternative scenarios with BEV fleets with mixed charging strategies.
3 Results
In the main part of the paper, we focus on the main scenario family where only part of the total BEV fleet considered in the reference switches to carsharing (“Shared + other BEVs”), i.e., the more policy-relevant scenario family. Results for the “Shared only” scenario family are provided in the appendix (D). Qualitative results are similar across scenario families unless explicitly mentioned.
3.1 Power sector cost increases remain moderate and are largest for bidirectional charging scenario
The adoption of carsharing in the BEV fleet results in increased power system costs across almost all scenario families, carsharing uptake regimes, and charging strategies (Figure 3 and Table 9). This cost increase is relatively small compared to the overall sector costs, and is always below 100 euros per substituted car per year in the main scenario family “Shared + other BEVs”, even in the case with bidirectional charging. In the “Shared + other BEVs” framework, it represents about 1.1% of the total sector costs and 2.2% in the “Shared only” framework. Note that the load of substituted BEVs is also small compared to the overall load and ranges between 1.1% and 2.6% of the overall load in the “Shared + other BEVs” scenario family, depending on the uptake regime.
Figure 3: Power sector costs by charging strategy and carsharing uptake (“Shared + other BEVs”)
In the case where all BEVs charge in an uncontrolled way, switching to electric carsharing implies very little additional cost or may even lead to slight cost reductions. This is because charging a large fleet of BEVs without considering power sector dynamics leads to a very expensive power sector, as BEVs considerably increase existing system peak loads (Figures 12 and 13). By switching to carsharing, the changing charging patterns may even slightly ease this burden. The fact that cost increases related to carsharing remain small should not be misinterpreted. The absolute sector costs remain the highest compared to other charging strategies, even after a switch to carsharing. Since the perspective of having a large BEV fleet charged in an uncontrolled way is neither realistic nor desirable (compare schill2015power (36)), we do not focus on this charging strategy in the rest of the paper. Results can nevertheless be found in the appendix (D).
Assuming that the whole BEV fleet charges smartly or bidirectionally, costs in the reference decrease compared to a setting with uncontrolled charging, as the former charging strategies help avoid costly load peaks and corresponding peaker plants and storage capacity (see section 3.2). In turn, switching to carsharing then increases costs, regardless of the uptake regime. This is because substituting private BEVs that are charged smartly or bidirectionally with shared BEVs partly takes away the flexibility that private BEVs provide to the power sector, as shared cars are driven more often. Note that we assume that private cars provide flexibility to the power sector with no additional investment costs. Substituting these private cars away implies an increase in the investment needed in the power sector. The cost increase is about twice as high when the whole BEV fleet is charged bidirectionally compared to when it is charged smartly, independent of the modeled uptake regime.
For both the smart and the bidirectional charging strategy, the increase in cost is higher when the carsharing uptake is higher. In the “Shared + other BEVs” scenario family, this holds not only in absolute terms, but also in relative terms: the cost increase in euros per substituted car is always greater, everything else being equal, when considering the high uptake. It ranges between 42 and 54 euros per substituted BEV in the smart charging case, and between 92 and 98 euros for bidirectional charging.
The greater the available flexibility options, the lower the cost increase due to a switch to carsharing. For instance, the increase in power sector cost is higher across all charging strategies and uptake rates when considering the “Shared only” framework, where no private BEV remain as a potential flexibility resource (Table 9). Private BEVs (which are absent in the “Shared only” scenario family) are more likely to be used flexibly since they are driven less and are thus available for grid interactions more often. The contrast between the two families is particularly clear-cut for the bidirectional charging strategy, where the cost increase per substituted car more than doubles in “Shared only” compared to “Shared + other BEVs”. The respective increase is only about 1.2 when assuming that all BEVs only charge smartly. This is because feeding back electricity into the grid is particularly useful in power sectors with high shares of renewables, as it helps to substitute peak generation and electricity storage capacities. Reducing the potential of a BEV fleet to engage in V2G accordingly hurts more than just reducing its potential to charge smartly.
Figure 4: Generation and storage discharging capacity by technology (“Shared + other BEVs”)
3.2 Changes in generation capacity and dispatch depend on charging strategy
The switch to carsharing triggers changes in the optimal electricity generation capacity mix that are different for smart and bidirectional charging, while the minimum renewable share constraint of 80% is binding. In the case of smart charging, the switch to carsharing reduces the potential to flexibilize the demand. In the case of bidirectional charging, switching to carsharing also entails a loss in the ability of the BEV fleet to feed back electricity into the grid when the demand is relatively high compared to the available supply. In all cases, the optimal PV generation capacity increases, by between around 2 and 7 GW (Figure 4, see also Figure 10).
In contrast, optimal wind power capacities generally decrease with carsharing. This is because the relatively large cumulative battery capacity of the fleet of privately owned BEV fosters the integration of wind power in the reference. This is visible in the cumulative BEV battery charging level time series, which not only shows diurnal patterns related to PV variability, but is dominated by longer-duration patterns that relate to wind power variability (Figures 24 and 25). If part of the BEV fleet switches to carsharing, the potential for such longer-duration storage decreases in line with the shrinking overall battery capacity. In turn, the share of wind power in the optimal capacity mix decreases, while PV with its shorter-duration diurnal variations increases. The only exception is the low uptake scenario with bidirectional charging, where generation from solar PV also decreases, and generation from bio energy increases (Figure 11). This is because the loss of bidirectional flexibility makes it favorable to use more dispatchable bio energy instead of a combination of more solar power and storage. Note that this switch from wind to solar power is not driven by the changes in BEV grid availability, as the latter is not the limiting factor even in the high uptake scenario (compare Figure 9).
Figure 5: Yearly electricity generation by technology (“Shared + other BEVs”)
When private BEVs with smart charging are substituted with shared BEVs, we observe an increase in optimal long-duration storage and PV capacities, and lower firm conventional capacity with high variable costs (OCGT) (Figure 4). As the fleet with carsharing is more frequently used for driving, it offers less room to shift the demand to extended periods with high availability of RES (e.g., July 6-8 in Figure 16).
In the case where all BEVs charge bidirectionally, the switch to carsharing entails additional investments into firm conventional capacity, in particular OCGT (about 1 to 3 additional GW) and to a lesser extent also CCGT (about 0.5 GW). Yet, this does not translate into a greater yearly electricity generation from gas power plants (Figure 5). The additional firm capacity is only needed during some periods in which the residual load is particularly high, and which are served by V2G of private BEVs in the reference, but to a lower extent by shared BEVs. In contrast to the smart charging case, the switch to carsharing does not increase electricity storage needs when BEVs charge bidirectionally. Here, the remaining fleet of private BEVs still provides such a high level of flexibility that there is no need for any investments in short- or long-duration storage (compare also Figures 22 and  23).
This, however, is not the case of in the “Shared only setting”, where no additional private BEVs are available. Hence, we observe additional investments into long-duration storage discharging capacity of almost 3 GW for both uptake regimes as well as into short-duration storage of about 5.5 GW in the low uptake regime (Figures 20 and 21). The same holds for storage charging capacity (Figures 18 and 19) and storage energy capacity (Figures 22 and  23). For the latter, carsharing increases the optimal long-duration storage energy capacity by around 300 GWh for both uptake regimes. This closely resembles the overall battery capacity of substituted private BEVs (Table 2). In the case with smart charging, the effects of carsharing on optimal electricity storage capacities are much smaller. This is because vehicle batteries do not provide as much flexibility to the electricity sector in this setting, and the flexibility loss triggered by carsharing accordingly requires less compensation in terms of electricity storage.
3.3 Shared BEVs can still charge and discharge in line with variable renewables, but to a lesser extent
We observe that, with perfect foresight, BEV fleets with shared cars could still be operated largely in line with variable renewable generation patterns (Figures 6 and 16). In particular, optimized charging takes place when RES availability is high, e.g., during negative residual load events in summer, and during periods with low positive residual load in winter (Figure 14). Conversely, BEVs are discharged (in the case of bidirectional charging) at times when the residual load is high, meaning that the demand exceeds variable renewable electricity generation.
A fleet with shared BEVs has a lower potential to balance renewable variations, i.e., to shift energy between negative and positive residual load periods, than a fleet of only private BEVs: light and dark blue areas on Figure 6 are always larger than or equal to the sum of the purple and pink areas. The difference in magnitude is smaller in the low uptake case (Figure 15). The charging and discharging patterns of the shared fleet (light purple and light pink) are more variable compared to those of the overall fleet. The reason is that shared vehicles are driven much more frequently and hence can only charge or discharge for shorter periods before they are used again. These results also apply to the “Shared only” setting (Figure 17).
Figure 6: Residual load curve, BEV charging and discharging for a subset of summer hours (upper panel) and winter hours (lower panel) [“Shared + other BEVs”, bidirectional charging, high uptake]. Note that the area plots are partly unstacked to better illustrate the decreasing flexibility of shared BEVs. The carsharing scenario areas (purple and pink) are stacked on top of each other, while the reference areas (blue) are not.
3.4 Green hydrogen production as another source of flexibility
In additional model runs, we include green hydrogen production as an example of another sector coupling technology which increases the demand-side flexibility of the system. More precisely, we consider a yearly industrial demand for green hydrogen of 30 TWh, which is assumed to be produced domestically via electrolysis. This largely reflects current German hydrogen policy targets. Note that this demand for green hydrogen comes on top of the hydrogen required for reconversion to electricity, i.e., for long-duration electricity storage. We assume that the industrial hydrogen demand is equally distributed over every hour of the year. In general, green hydrogen can be flexibly produced at low costs in periods when plenty of renewable electricity is available and prices are low. Yet, shifting hydrogen production to such periods requires according electrolysis and hydrogen storage capacities. The model optimizes these capacities as well as their hourly use as endogenous model variables, rendering hydrogen production a flexible sector coupling option.
Including additional green hydrogen production in the model leads to lower overall cost increases of carsharing both in absolute and relative terms (Table 9). Additional costs per substituted car decrease by 14-22% in the “Shared + other BEVs” framework and by 18-32% in the “Shared only” framework, compared to the setting with no green hydrogen industrial demand. The reason for this is that green hydrogen production competes with smartly charged BEVs for cheap electricity in periods where renewable availability is high. Accordingly, smart and bidirectional charging are less beneficial in the reference than in the setting without hydrogen. In turn, reducing the BEV fleet’s flexibility potential by shifting to carsharing also hurts less.
This is also visible in the effects of carsharing on optimal generation and storage capacities. Irrespective of the uptake scenario, carsharing triggers lower overall investments into renewable and firm generation capacity as well as electricity storage as compared to the case without green hydrogen (Figures 26 to 29). In particular, optimal renewable capacities decrease. Besides, the reduction in wind generation capacity is even more pronounced in the case with a green hydrogen industrial demand. This is because the optimal renewable share exceeds 80% in the reference with industrial green hydrogen demand.
4 Discussion of model choices and limitations
Our model results rely on a set of assumptions. The parameterization is on purpose relatively restrictive when it comes to the flexibility of the power sector: it does not include interconnection with other neighbouring countries, nor other flexible demand-side options that could engage in temporal load shifting. Including additional flexibility options would likely decrease the impacts of carsharing, which would translate into smaller cost increases as well as smaller deviations from the reference in terms of capacity mix and dispatch. In this respect, our main results can be interpreted as an upper bound of the power sector impacts.
When it comes to the specific mobility time series, our framework is also rather conservative in terms of flexibility. In the “Shared only” framework, we consider only one cluster that switches to carsharing, namely the one with the smallest amount of trips per day and with the shortest overall travelled distance. These vehicles are idle most of the time, and thus offer the greatest flexibility potential if smart charging or vehicle-to-grid technologies are available. Assuming that such profiles switch entirely to carsharing removes the greatest source of flexibility that one could get from the sector coupling with passenger cars. The results thus provide a very conservative quantification of how much the flexibility potential of BEVs would be reduced. The same applies to the “Shared + other BEVs” framework, although to a smaller extent, since we also consider most intense car usage by including clusters 3 to 5. Cost increases still remain moderate.
Another key assumption of our setting is that car mobility needs remain constant, irrespective of the way they are fulfilled. This means that the switch to carsharing does not induce a modification of when and how people decide to move with a car. This is disputable as it could be the case that switching to carsharing comes along with less car trips or shorter car trips, potentially combined with other transport modes such as active modes or public transport. In the case where carsharing would induce a decrease in overall car use, this could entail a greater flexibility of the shared cars than the one we model in this paper, for a constant size of the shared BEV fleet. This corroborates our upper bound interpretation of the modeled flexibility reduction effects.
Next, car users and fleet operators are assumed to have perfect foresight when it comes to charging and discharging their cars. This means that people can perfectly predict their car trips over time. This absence of uncertainty when it comes to the timing and duration of future trips enables to optimize the grid interactions of the overall fleet to the fullest extent possible, providing a benchmark that may not be achievable with more realistic behaviours. In that sense, our modelling might overestimate the flexibility of both private and shared BEVs. Note that we focus on deviations to respective references, such that assumptions present in both settings play less of a role.
Some modelling features rely on stylized assumptions for tractability and numerical feasibility, especially concerning BEVs time series. For example, including more car mobility profiles and additional vehicle models would be desirable to generate an even more representative fleet, but this heavily increases the numerical burden. Likewise, we assume that shared cars always find a charging station after a trip and that such charging stations have a power rating of 75 kW. Even if this is optimistic compared to today’s situation, it appears plausible that future carsharing fleets could be managed by their operators such that users can always find a charging station, potentially a specifically designed one for carsharing users.
On the power modelling side, a potential limitation is the optimization at a national scale and not at the substation level, which could potentially overestimate the flexibility potential of BEVs with smart charging (strobel2022joint (38)).
5 Conclusion
The results indicate that a switch to electric carsharing decreases the flexibility potential of BEVs and thus increases electricity sector costs. This cost increase however remains moderate in most settings, and remains below 100 euros per substituted car and per year in our main scenario. Cost increases are highest when other sources of flexibility are not available, such as private BEVs or green hydrogen production, and when all electric vehicles are assumed to be charged bidirectionally, i.e., make full use of their flexibility potential.
Costs effects of carsharing are driven by changes in the optimal power plant portfolio and its hourly use, which also sheds light on the benefits of coupling a largely renewable power sector with BEVs. In general, switching to carsharing increases the optimal share of solar PV and in turn decreases the share of wind power, as the aggregate vehicle battery storage capacity decreases. Effects on stationary electricity storage needs depend on various factors. In our main scenario, we find that the increase in stationary storage caused by carsharing is much lower than the substituted battery capacity of privately owned cars. Concerning the use of mineral resources for batteries, both for vehicles and stationary storage, carsharing is thus likely to have an overall relieving effect.
Overall, our research demonstrates that future BEV fleets that include relevant amounts of shared cars can still be operated largely in line with variable renewable energy sources and thus contribute to their system integration. Note that this relies on the assumption that vehicle operators, or some aggregators, have perfect foresight on mobility patterns and charging availability, both for privately-owned and shared BEVs. Hence, our results imply an upper bound of the beneficial system effects of flexible BEVs. Yet, we argue that system-oriented charging of future shared car fleets appears to be plausible, given that carsharing operators are likely to have an interest in low charging costs.
While most effects determined in our analysis are relatively small in size, both in terms of cost increases and changes in the capacity mix, they should be put in perspective with potential gains from switching to electric carsharing and reducing the BEV fleet size. BEVs are also associated with negative externalities which can be thought as social costs (gossling2019social (15, 7)). Methods to derive estimates of such costs have been developed (schroder2023ending (37)). In a case study for Munich, the overall external cost of a BEV is estimated to be around 14 euro-cents per kilometer driven. Assuming BEVs are driven ten kilometers a day, and five days per week, the yearly external cost of a BEV is around 364 euros. Removing one BEV from the system could thus represent a gain of 364 euros per year for society. In our main scenario, these reduced negative externalities are around four times larger than power sector cost increases under the assumption of bidirectional charging, and even around a factor eight larger assuming smart charging. These estimates need to be interpreted cautiously since they depend on many underlying assumptions and are specific to each urban setting. Nevertheless, they highlight that the increase in the power sector costs is only a part of the overall impact of switching from private to shared cars. In addition, carsharing may also reduce the carbon footprint of BEVs because of higher overall use of cars over their lifetime (morfeldt2022 (30)). We conclude that the potential societal gains from the reduction of the car fleet size are likely to outweigh the potentially adverse electricity sector effects. In this light, incentives and support to implement effective and profitable sharing services appear meaningful along with further research on how to most optimally design and operate such fleets and the charging infrastructure supporting them.
Author contributions
Adeline Guéret: Conceptualization, Methodology, Software, Formal analysis, Investigation, Data Curation, Writing - Original Draft, Writing - Review & Editing, Visualization. Wolf-Peter Schill: Conceptualization, Investigation, Writing - Review & Editing, Funding acquisition. Carlos Gaete-Morales: Software.
Acknowledgments
[…]
Appendix A Clusters
Figure 7: Dendrograms for the sequences for all location types (working days)
Metropolises Big cities Middle-size cities Small cities Rural areas
Number of sequences
Cluster 1 107 63 143 67 130
Cluster 2 363 158 194 298 656
Cluster 3 790 561 908 1,877 1,777
Cluster 4 2,984 2,072 2,277 6,437 2,931
Cluster 5 3,188 3,491 6,451 5,789 7,156
Cluster 6 3,662 5,513 9,504 11,725 10,664
Total 11,094 11,858 19,477 26,193 23,314
Shares of clusters in each location
Cluster 1 1.0 0.5 0.7 0.3 0.6
Cluster 2 3.3 1.3 1.0 1.1 2.8
Cluster 3 7.1 4.7 4.7 7.2 7.6
Cluster 4 26.9 17.5 11.7 24.6 12.6
Cluster 5 28.7 29.4 33.1 22.1 30.7
Cluster 6 33.0 46.5 48.8 44.8 45.7
Total 100 100 100 100 100
Average number of trips per day
Cluster 1 2.3 2.1 2.2 2.0 2.2
Cluster 2 3.5 2.8 3.0 2.9 3.5
Cluster 3 3.9 3.9 3.9 3.9 4.2
Cluster 4 3.6 3.9 4.0 3.9 3.7
Cluster 5 2.8 3.4 3.6 3.2 3.4
Cluster 6 2.1 2.3 2.5 2.5 2.5
Overall daily travelled distance (in kilometers)
Cluster 1 410.3 407.2 340.6 328.5 276.3
Cluster 2 239.3 360.5 310.3 324.6 243.2
Cluster 3 112.6 156.4 160.4 150.0 123.3
Cluster 4 49.7 68.3 82.2 63.4 74.2
Cluster 5 25.2 32.9 39.5 34.7 46.6
Cluster 6 10.6 13.4 13.2 14.0 17.1
Average trip distance (in kilometers)
Cluster 1 249.2 256.2 228.5 215.2 172.9
Cluster 2 116.1 186.9 162.2 165.4 108.5
Cluster 3 43.8 65.2 65.1 60.2 41.2
Cluster 4 18.5 24.4 28.7 21.5 26.5
Cluster 5 11.1 12.5 14.2 13.4 16.8
Cluster 6 5.5 6.2 6.0 6.4 7.9
Average trip duration (in minutes)
Cluster 1 343.7 438.7 375.0 485.6 421.6
Cluster 2 126.6 189.1 164.8 214.5 127.5
Cluster 3 57.9 79.2 79.0 73.5 52.5
Cluster 4 34.2 37.8 41.1 32.6 36.7
Cluster 5 24.5 24.3 25.0 22.8 24.8
Cluster 6 14.1 14.5 13.0 12.5 13.0
Table 3: Cluster statistics by location (working days)
Figure 8: Sequence index plots (metropolises, working days, clusters 1 to 6)
Appendix B emobpy
emobpy is an open-source Python tool designed to generate synthetic BEV time series. Using the tool, four different types of time series can be generated: (i) vehicle mobility, (ii) driving electricity consumption, (iii) grid availability and (iv) grid electricity demand. This order matters since each time series builds upon the previous one in order to be created.
The vehicle mobility time series indicates how many trips are undertaken during a day and, for each trip the destination, distance and duration. Possible destinations are work/school, leisure, home and errands. The generation of the vehicle mobility time series for shared BEVs relies on the same distributions as for private BEVs of a given cluster in a given location but centered around a new mean. To calculate the mean number of trips for shared BEVs, we first calculate the total number of private cars in a cluster, assuming that each sequence corresponds to one car. We then compute the total number of shared cars based on the assumed substitution rate. Finally, we calculate the average number of trips per shared car in the cluster.
The driving electricity consumption time series computes for 5-minutes time step how much electricity is used to power the trip depending on vehicle and climatic features such as the motor or the ambient temperature as well as the speed of the trip.
The grid availability time series is also computed at the 5-minutes time step and indicates whether the vehicle is plugged to a charging station and if yes, at which power rating.
Finally, the grid electricity time series corresponds to the electricity actually taken from the grid for charging a vehicle, using the consumption and availability time series and assuming a given charging behaviour. We only compute this last time series when looking at scenarios with uncontrolled charging.
Table 4: Conditionality criteria and levels for empirical probability distributions
Weekday (Mo-Fri) Saturday Sunday
Conditionality Level Conditionality Level Conditionality Level
Number of trips no no no
Destination departure time
number of trips
trip rank
number of trips
trip rank
number of trips
trip rank
Private BEVs Joint distance/duration
number of trips
destination
cluster x location
number of trips
destination
location
number of trips
destination
location
Number of trips no no no
Destination departure time no no no
Shared BEVs Joint distance/duration destination cluster x location destination location destination location
Table 5: emobpy assumptions for generating grid availability time series
Charging station Power rating Battery capacity
destination availability 0 kW 3.7 kW 11 kW 22 kW 75kW 150 kW (kWh)
Private BEVs home 0.9 0.1 0.6 0 0.3 0 0 58
work/school 0.9 0.1 0 0.3 0.5 0.1 0
errands 0.8 0.2 0 0.2 0.5 0.1 0
leisure 0.8 0.2 0 0.2 0.5 0.1 0
driving 0.01 0.99 0 0 0 0.005 0.005
Shared BEVs home 1 0 0 0 0 1 0 100
work/school 1 0 0 0 0 1 0
errands 1 0 0 0 0 1 0
leisure 1 0 0 0 0 1 0
driving 0.01 0.99 0 0 0 0.005 0.005
Figure 9: Battery electric vehicles’ average overall grid availability for each hour of the week in the reference and carsharing scenarios [“Shared + other BEVs”]
Appendix C DIETER
Metropolises Big cities Middle-sized cities Small-sized cities Rural areas Total
Cars (%) 11 12 22 27 29 100
Cars (million units) 1.62 1.73 3.30 4.05 4.29 15
Table 6: Distribution of cars across location types
Capacity bounds
Lower Upper Interest rates Lifetime Availability Overnight costs Fixed costs Efficiency Carbon content Fuel costs
[in GW] [in GW] [years] [1,000 EUR/MW] [1,000 EUR/MW] [t/MWh] [EUR/MWh]
Hard coal 0.0 9.8 0.04 35 0.955 3845 30 0.430 0.337 8.27
Lignite 0.0 9.3 0.04 35 0.953 3845 30 0.380 0.399 5.50
Natural gas (OCGT) 0.0 +Inf 0.04 25 0.948 435 15 0.400 0.201 30.00
Natural gas (CCGT) 0.0 +Inf 0.04 25 0.960 830 20 0.542 0.201 30.00
Bio energy 6.0 6.0 0.04 30 1.0 3300 100 0.487 0.000 32.50
Run-of-river 5.6 5.6 0.04 50 1.0 3312 30 0.900 0.000 0.00
Photovoltaic 0.0 +Inf 0.04 25 1.0 519 25 1.000 0.000 0.00
Onshore wind 0.0 +Inf 0.04 25 1.0 1035 35 1.000 0.000 0.00
Offshore wind 0.0 +Inf 0.04 25 1.0 1934 100 1.000 0.000 0.00
Table 7: Cost and technology parameters for electricity generation technologies
Capacity bounds Overnight costs Efficiency Marginal costs
energy power in/out Interest rate Lifetime Availability energy charging power discharging power charging discharging charging dicharging
lower upper lower upper
[GWh] [GW] [years] [1,000 EUR/MWh] [1,000 EUR/MW] [1,000 EUR/MW] [EUR/MW] [EUR/MW]
Lithium-ion batteries 0 +Inf 0 +Inf 0.04 20 0.98 142 80 80 0.96 0.96 0.5 0.5
Long-duration storage
Electrolysers 0 +Inf 0 +Inf 0.04 10 1.0 - 40 - 0.71 - 0.5 -
Compressor 0 +Inf 0 +Inf 0.04 15 1.0 - 80 - 0.995 - 0.5 -
Cavern storage 0 +Inf 0 +Inf 0.04 100 0.95 2 - - - - - -
Reconversion 0 +Inf 0 +Inf 0.04 25 0.95 - - 435 - 0.4 - 0.5
Pumped-hydro storage 57 57 9.5 9.5 0.04 80 0.98 10 550 550 0.97 0.91 0.5 0.5
Table 8: Cost and technology parameters for electricity storage technologies
Appendix D Detailed results for all scenarios
D.1 Power sector costs
Shared only
Uncontrolled Smart charging Bidirectional
Low High Low High Low High
Reference (in billion euros) 56.3 57.6 55.6 56.2 54.4 54.8
Difference (in million euros) 21.6 8.2 196.5 429.4 977.5 1214.0
[…]
Reference (in billion euros) 58.3 59.6 57.7 58.4 56.7 57.3
Difference (in million euros) 19.7 14.3 160.4 295.4 734.3 847.6
[…]
Shared + other BEVs
Uncontrolled Smart charging Bidirectional
Low High Low High Low High
Reference (in billion euros) 61.5 61.5 58.4 58.4 57.0 57.0
Difference (in million euros) 6.6 -15.2 163.4 338.3 355.0 613.3
[…]
Reference (in billion euros) 63.5 63.5 60.9 60.9 59.7 59.7
Difference (in million euros) 5.3 -12.1 132.7 263.9 302.5 518.9
[…]
Table 9: Power sector costs for reference and scenarios across frameworks, uptake regimes and charging strategies
D.2 Power generation and capacity mix
Figure 10: Effects of carsharing on optimal generation and storage discharging capacity
Figure 11: Effects of carsharing on yearly electricity generation
D.3 Power dispatch
Figure 12: Residual load for different charging strategies for a subset of summer hours [“Shared + other BEVs”, reference]
Figure 13: Residual load for different charging strategies for a subset of winter hours [“Shared + other BEVs”, reference]
Figure 14: Hourly dispatch for a subset of summer hours (upper panel) and winter hours (lower panel) [“Shared + other BEVs”, bidirectional charging, high uptake]
Figure 15: Residual load curve, BEV charging and discharging for a subset of summer hours (upper panel) and winter hours (lower panel) [“Shared + other BEVs”, bidirectional charging, low uptake]
Figure 16: Residual load curve, BEV charging and discharging for a subset of summer hours (upper panel) and winter hours (lower panel) [“Shared + other BEVs”, smart charging, high uptake]
Figure 17: Residual load curve, BEV charging and discharging for a subset of summer hours (upper panel) and winter hours (lower panel) [“Shared only”, bidirectional charging, high uptake]
D.4 Storage capacity and dispatch
Figure 18: Storage charging power capacity in reference cases
Figure 19: Effects of carsharing on optimal storage charging power capacity
Figure 20: Storage discharging power capacity in reference cases
Figure 21: Effects of carsharing on optimal storage discharging power capacity
Figure 22: Storage energy capacity in reference cases
Figure 23: Effects of carsharing on optimal storage energy capacity
Figure 24: Battery electric vehicles’ overall storage level in the reference and the corresponding carsharing scenario [“Shared + other BEVs”, smart charging, high uptake]
Figure 25: Battery electric vehicles’ overall storage level in the reference and the corresponding carsharing scenario, zoomed in for a subset of winter hours [“Shared + other BEVs”, smart charging, high uptake]
D.5 Industrial hydrogen demand
Figure 26: Power generation capacity mixes in reference settings comparing without and with industrial demand for green hydrogen [“Shared + other BEVs”]
Figure 27: Power generation capacity mixes in carsharing scenarios comparing without and with industrial demand for green hydrogen [“Shared + other BEVs”]
Figure 28: Storage energy capacity in reference settings comparing without and with industrial demand for green hydrogen [“Shared + other BEVs”]
Figure 29: Storage energy capacity in carsharing scenarios comparing without and with industrial demand for green hydrogen [“Shared + other BEVs”]